# DiMoDif: Discourse Modality-information Differentiation for Audio-visual Deepfake Detection and Localization

### v1

@@ Abstract
Deepfake technology has rapidly advanced, posing significant threats to information integrity and societal trust. While significant progress has been made in detecting deepfakes, the simultaneous manipulation of audio and visual modalities, sometimes at small parts but still altering the meaning, presents a more challenging detection scenario. We present a novel audio-visual deepfake detection framework that leverages the inter-modality differences in machine perception of speech, based on the assumption that in real samples – in contrast to deepfakes – visual and audio signals coincide in terms of information. Our framework leverages features from deep networks that specialize in video and audio speech recognition to spot frame-level cross-modal incongruities, and in that way to temporally localize the deepfake forgery. To this end, DiMoDif employs a Transformer encoder-based architecture with a feature pyramid scheme and local attention, and optimizes the detection model through a composite loss function accounting for frame-level detections and fake intervals localization. DiMoDif outperforms the state-of-the-art on the Temporal Forgery Localization task by +47.88% AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF. On the Deepfake Detection task, it outperforms the state-of-the-art by +30.5% AUC on AV-Deepfake1M, +2.8% AUC on FakeAVCeleb, and performs on-par on LAV-DF. Code available at https://github.com/mever-team/dimodif.
@@ 1 Introduction
Audio-visual deepfakes are AI-generated content involving manipulations to one or both modalities, sometimes at small time intervals, with the intention to deceive [45]. Deepfake content can widely spread online, and mislead viewers, contributing to the growing issue of information disorder. A common feature of audio-visual deepfakes is the presence of incongruities between the visual and audio signals, which are, however, increasingly hard to detect manually. This necessitates the development of robust AI-based deepfake detection methods.
Deepfake detection is a growing area of research [54, 51, 46]. Existing approaches primarily focus on pixel-level analysis, which examines inconsistencies in gradient variations [65], color variations [30], or artifacts introduced during the generation process [11]. Also, feature-based methods analyze facial landmarks [72], temporal coherence [27], and other visual cues to identify inconsistencies in deepfake content [40]. Furthermore, there has been growing interest in detecting and localizing deepfakes in audio-visual content, with both tasks being particularly challenging due to the complex interplay between the two modalities. Key approaches include audio-visual synchronization that analyzes the consistency between visual and auditory cues namely face movements and speech [26, 9, 14], feature reconstruction learning with cross-reconstruction similarity estimation to detect anomalies prevalent in fake videos [76], and self-supervised approaches relying on the synchronization patterns learned from real videos only [28, 18].
@@
(a) Example of speech recognition differences between lip-reading and speech-to-text models.
(b) Aggr. FakeAVCeleb [34].
(c) Aggr. LAV-DF [10].
(d) Aggr. AV-Deepfake1M [8].
Figure 1: The process of identifying audio-visual inconsistency for deepfake detection. In (a), a video’s visual and audio components are separately processed by lip-reading and speech-to-text models, then a difference score is calculated. In (b,c,d) the difference score distributions are illustrated in the form of violin plots for three different datasets.
@@
We illustrate how state-of-the-art AI analysis tools could emulate this process in Figure 1. In Figure 1(a), an input video is first decomposed into its constituent visual and audio streams. These are then fed into state-of-the-art models for lip-reading and speech-recognition [44], and the resulting textual outputs are compared [^1] Python’s $\mathtt{difflib}$ is used (cf. https://docs.python.org/3/library/difflib.html).. Their difference is computed using the normalized Levenshtein distance, i.e. the ratio between the edit distance $\mathtt{delta}$ of the two sentences and the sum of their lengths, which results in scores between 0% (identical) and 100% (completely different). Figures 1(b), 1(c) and 1(d) illustrate aggregate results through violin plots computed on the evaluation sets of three popular audio-visual deepfake detection benchmarks, FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8]. On average, lower difference is observed in real videos compared to fake ones across all datasets, indicating the suitability of this score as a measure of incongruity between audio and video. Even though these differences are significant on FakeAVCeleb, they are smaller on LAV-DF, and much smaller on AV-Deepfake1M, which render naive thresholding on this score unsuitable for robust detection. This is due to the fact that FakeAVCeleb’s videos are fully manipulated, while LAV-DF’s samples are manipulated only in parts, while samples from AV-Deepfake1M are manipulated in even smaller parts (on average half the length of those in LAV-DF).
@@
Motivated by the above, we propose Discourse Modality-information Differentiation (DiMoDif), a deep learning-based deepfake detection architecture that exploits the inter-modality differences in machine perception of speech. DiMoDif first decomposes the input video into its visual and audio streams, and extracts the respective video and audio representations based on state-of-the-art lip reading and speech-to-text pre-trained models [44]. At its core, DiMoDif employs a Transformer encoder architecture [66] with a feature pyramid scheme and local attention configuration, to spot frame-level cross-modal incongruities between the two modality-specific representations, and in that way detect and localize the deepfake forgeries. Note that low difference scores frequently occur in partly manipulated fake videos (e.g. Figure 1(a)); however, DiMoDif manages to classify them as fake by capturing audio-visual divergence at the frame level. In addition, high difference scores occasionally occur in real videos; however, DiMoDif leverages alternative features, e.g., hard phonemes [6], to classify them as real (cf. Section 5.2). The architecture is optimized using a composite loss function that accounts for frame-specific detections, overlaps between predicted and ground truth fake intervals, and divergence in corresponding boundaries. We evaluate DiMoDif on Deepfake detection (DFD) and Temporal Forgery Localization (TFL) tasks using three audio-visual deepfake detection benchmarks, FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8]. DiMoDif outperforms the state-of-the-art on FakeAVCeleb by +2.8% AUC, exhibits a significant performance increase of +30.5% AUC and +47.88% AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF dataset. We also provide an analysis of DiMoDif’s generalization abilities, and an extensive ablation analysis to ground our methodological choices.
@@ 2.1 Deepfake Detection Datasets
During the past years, several datasets have been proposed for video deepfake detection. Most of them focus solely on visual manipulations, thus including only the visual modality in their samples, e.g., DF-TIMIT [35], FaceForensics++ [58], DeeperForensics [32], Celeb-DF [41], WildDeepfake [81], and DF-Platter [48]. Our focus is on the detection of audio-visual deepfakes, which require content, manipulations and annotations in both modalities. Table 1 presents deepfake detection datasets that provide both visual and audio streams. DFDC [20] is a seminal large-scale benchmark in the field that contains approximately 128K samples. Although it provides audio-visual content and contains manipulations on both modalities, (i) it does not provide separate manipulation annotations per modality, (ii) fake audio parts appear in only 4% of the total number of videos, while fake visual parts in 84% [^2] Based on a later estimation: https://www.kaggle.com/datasets/basharallabadi/dfdc-video-audio-labels., and (iii) it contains videos where the depicted individual does not talk or the speaking person is not depicted and the depicted one does not talk. These features render DFDC inappropriate for audio-visual deepfake detection. KoDF [36] is another large-scale benchmark containing 237K videos of Korean subjects. Although it provides both the visual and audio components of its samples, it contains only visual manipulations, which makes it unsuitable for our problem. FakeAVCeleb [34], contains 21K manipulated and 0.5K real videos with manipulations on both modalities and corresponding annotations. Due to its size it is commonly used as an evaluation benchmark; here, we consider it for training and evaluation purposes with real video additions from the VoxCeleb2 dataset [15]. LAV-DF [10] is a recently introduced dataset that contains 136K real and fake videos with manipulations and targets on both modalities for two tasks, Deepfake Detection (DFD) and Temporal Forgery Localization (TFL). Finally, AV-Deepfake1M [8] is the largest audio-visual deepfake detection benchmark to date containing over 1.1M real and fake videos with manipulations and targets on both modalities, for both the DFD and TFL tasks. In our experimental analysis, we consider FakeAVCeleb, LAV-DF, and AV-Deepfake1M datasets, which meet all requirements of the audio-visual deepfake detection task. We also consider DFDC and KoDF for cross-dataset generalization evaluation experiments.
@@
2.2 Deepfake Detection Approaches
Many deepfake detection approaches consider only video-level manipulations and processing [62, 58, 78, 38, 1, 17, 55]. Although these are only partly related to DiMoDif  we compare our method with them following previous practice. Recent works in deepfake detection have focused on leveraging both visual and audio information through learning mechanisms based on multimodal fusion, contrastive learning, and self-supervised-learning, aiming to uncover cross-modal inconsistencies in deepfake videos. Specifically, [80] demonstrated the effectiveness of an attention-based detection framework that identifies the violation of cross-modal synchronization patterns in deepfake videos, while [71] proposed a framework exploiting audio-visual inconsistency by encoding spatio-temporal information, fusing multimodal features, and then classifying based on inter- and intra-modal disharmony levels. Similarly, [56, 31, 82] propose joint representation learning frameworks that consider inter- and intra-modal encoding mechanisms. UMMAFormer [76] adopts feature reconstruction and cross-reconstruction attention combined with an enhanced feature pyramid network, while BA-TFD & BA-TFD+ [10, 9] adopt 3DCNN and multi-scale vision Transformer backbones guided by contrastive, frame classification, and boundary matching objectives. The contrastive learning approach is also adopted by other works for comparing inter-modal representations within a video [14] or between different videos [77]. Other papers focus on feature similarity to identify audio-visual inconsistencies [26], and [59] compares features of the video’s lip sequence with a lip sequence generated from the audio input. Self-supervised approaches that leverage intrinsic synchronization of real videos have also been utilized [75, 25, 28, 22, 18, 53, 73, 21] to tackle the generalization challenge faced by methods trained on specific deepfake generation models. A similar motivation with ours is presented in [7, 39], which rely on a simple threshold on the divergence between visual and audio speech recognition model outputs. However, as shown in Figure 1 this simple technique is insufficient on benchmarks with harder samples than FakeAVCeleb, such as LAV-DF and AV-Deepfake1M. In contrast, our approach extracts lip-reading and speech-to-text content representations that are then processed by a Transformer-based fusion mechanism that is capable of identifying discriminative speech features on top of speech divergence ones. Pre-trained lip-reading models have also been utilized as feature extractors in [29], but for visual-only deepfake detection. In addition, [57] sets a threshold on cross-modal feature similarity deriving from an audio-visual speech self-supervised representation learning model. Also, specialized representation learning has been explored by [47] that considers modality emotion consistency and [69] that considers articulatory and lip movement consistency. Finally, several works [42, 50, 61, 74, 5] of the Temporal Action Localization (TAL) domain are considered as relevant by the deepfake detection literature.
@@ 3.1 Problem Formulation
Consider a dataset $\mathfrak{D}=\{\mathfrak{v}_{i},y_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ where $\mathfrak{v}_{i}=\{v,a\}$ denotes a video, containing the visual $v\in\mathbb{R}^{h\times w\times c\times f}$ and the audio $a\in\mathbb{R}^{s}$ signal [^3] Utterances mainly comprise monophonic signals, besides conversion to single-channel is straightforward if required., with $h$ denoting the height, $w$ the width, $c$ the number of channels, $f$ the number of video frames, and $s$ the number of audio samples. $y_{i}\in\{0,1\}^{2}$ denotes the Deepfake Detection (DFD) target per modality, while $\mathbf{y}_{i}=\{y^{v}_{i,1},\dots,y^{v}_{i,f},y^{a}_{i,1},\dots,y^{a}_{i,f}\}$, with $y^{m}_{i,j}=(d^{m,s}_{i,j},d^{m,e}_{i,j},a^{m}_{i,j})$, denotes the Temporal Forgery Localization (TFL) target, assigning a forgery ground truth $a^{m}_{i,j}\in\{0,1\}$ to each modality $m\in\{v,a\}$ and frame $j$, along with the corresponding distances $d^{m,s}_{i,j},d^{m,e}_{i,j}\in\mathbb{R}$ between the frame $j$ and the onset $s^{m}_{i,j}$ and offset $e^{m}_{i,j}$ of the fake interval, if the frame $j$ is fake. We train deep networks $\mathcal{F}$, the architecture of which is analysed in Section 3.2, on the DFD task $\hat{y}_{i}=\mathcal{F}(\mathfrak{v}_{i})$ and the TFL task $\hat{\mathbf{y}}_{i}=\mathcal{F}(\mathfrak{v}_{i})$ separately, while the predictions $\hat{\mathbf{y}}_{i}$ are decoded by:
$\hat{a}^{m}_{i,j}>0.5\text{, }\hat{s}^{m}_{i,j}=j-\hat{d}^{m,s}_{i,j}\text{, }\hat{e}^{m}_{i,j}=j-\hat{d}^{m,e}_{i,j}$ (1)
$\hat{\cdot}$ denotes prediction while its absence ground truth. Henceforth, we omit sample index $i$ to reduce notation complexity.
3.2 Architecture

### v2

@@ Abstract
Deepfake technology has rapidly advanced and poses significant threats to information integrity and trust in online multimedia. While significant progress has been made in detecting deepfakes, the simultaneous manipulation of audio and visual modalities, sometimes at small parts or in subtle ways, presents highly challenging detection scenarios. To address these challenges, we present DiMoDif (fig. 1), an audio-visual deepfake detection framework that leverages the inter-modality differences in machine perception of speech, based on the assumption that in real samples – in contrast to deepfakes – visual and audio signals coincide in terms of information. DiMoDif leverages features from deep networks that specialize in visual and audio speech recognition to spot frame-level cross-modal incongruities, and in that way to temporally localize the deepfake forgery. To this end, we devise a hierarchical cross-modal fusion network, integrating adaptive temporal alignment modules and a learned discrepancy mapping layer to explicitly model the subtle differences between visual and audio representations. Then, the detection model is optimized through a composite loss function accounting for frame-level detections and fake intervals localization. DiMoDif outperforms the state-of-the-art on the Deepfake Detection task by 30.5 AUC on the highly challenging AV-Deepfake1M, while it performs exceptionally on FakeAVCeleb and LAV-DF. On the Temporal Forgery Localization task, it outperforms the state-of-the-art by 47.88 AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF. Code available at https://github.com/mever-team/dimodif.
@@ 1 Introduction
Figure 1: Partial audio-visual manipulation leads to different visual and audio speech predictions. DiMoDif detects and localizes the fake part based on feature space incongruity.
Audio-visual deepfakes are AI-generated multimedia involving manipulations to one or both modalities, sometimes in small parts, with the intention to deceive [47]. Deepfake content can spread widely online, and mislead viewers, contributing to the growing issue of information disorder. A revealing feature of audio-visual deepfakes is the presence of incongruities between the visual and audio signals, which are, however, increasingly hard to detect due to the quality and realism of current deepfakes. This necessitates the development of robust AI-based deepfake detection methods.
Deepfake detection is a growing research area [55]. Pertinent methods focus on pixel-level (gradient [69], color [31], and artifact analysis [11]) and feature-based (facial landmarks [76], temporal coherence [28]) inconsistencies. An expanding area of study involves the detection and localization of deepfakes in audio-visual data, which is especially hard due to the way sound and visuals are intertwined. Key approaches analyze the consistency between raw visual and auditory cues [27, 9, 15], consider feature reconstruction learning modules [80], and follow self-supervised approaches learning synchronization patterns from real videos only [29, 19].
(a) Example of differences between Visual Speech Recognition (VSR) and Audio Speech Recognition (ASR) outputs.
File path: AV-Deepfake1M/val/id01003/9nmM17wsxGU/00013/fake_video_fake_audio.mp4
Figure 2: Identifying machine perception discrepancies between visual and audio speech for deepfake detection. In (a), a video’s visual and audio streams are separately processed by VSR and ASR models, then the outputs’ normalized Levenshtein distance $\mathtt{d_{L}}$ is calculated. In (b,c,d) the $\mathtt{d_{L}}$ distributions are illustrated for FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8].
@@
We illustrate how state-of-the-art AI analysis tools could emulate this process in fig. 2. In fig. 2(a), an input video is decomposed into its visual and audio streams feeding state-of-the-art models for Visual and Audio Speech Recognition (VSR, ASR) [^1] Also known as lip-reading and speech-to-text, respectively. [46]. The difference between the two textual outputs is quantified using the normalized Levenshtein distance ($\mathtt{d_{L}}$), calculated by dividing the edit distance (delta) [^2] Determined through Python’s difflib. by the combined length of the texts, resulting in a percentage from 0% (identical) to 100% (completely different). Figures 2(b), 2(c) and 2(d) illustrate aggregate results computed on the evaluation sets of three popular audio-visual deepfake detection benchmarks, FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8]. Real videos, on average, showed lower difference scores $\mathtt{d_{L}}$ than fake videos across all datasets, confirming the score’s ability to measure incongruity. However, even though these differences are significant on FakeAVCeleb, they are smaller on LAV-DF, and much smaller on AV-Deepfake1M, which renders naive thresholding unsuitable for robust detection. This is due to the fact that FakeAVCeleb’s videos are fully manipulated, while LAV-DF’s samples are manipulated only in parts, and samples from AV-Deepfake1M are manipulated in even smaller parts (on average half the length of those in LAV-DF).
@@
To leverage machine perception discrepancies between visual and audio speech, we introduce Discourse Modality-information Differentiation (DiMoDif). The proposed architecture decomposes the input video into its visual and audio streams, and extracts features via the Discourse-related Feature Extraction (DiFE) module, which uses the pre-trained VSR and ASR models proposed in [46]. Subsequently, the Modality-information Differentiation (MiD) module, implemented via a Transformer encoder with local cross-modal attention and feature pyramids, identifies frame-level inconsistencies and utilizes them for deepfake detection and localization. Note that low difference scores $\mathtt{d_{L}}$ frequently occur in partly manipulated fake videos (e.g. fig. 2(a)); however, DiMoDif manages to classify them as fake by capturing audio-visual divergence at the frame level. In addition, high difference scores $\mathtt{d_{L}}$ occasionally occur in real videos; however, DiMoDif leverages alternative features, e.g., hard phonemes [6], to classify them as real (cf. section 5.2). A composite loss optimizes frame-level detection and fake interval prediction (overlap and boundaries). We evaluate DiMoDif on Deepfake detection (DFD) and Temporal Forgery Localization (TFL) tasks using three audio-visual deepfake detection benchmarks, FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8]. DiMoDif outperforms the state-of-the-art by a considerable margin on the challenging AV-Deepfake1M (+30.5% AUC for DFD and +47.88% AP@0.75 for TFL), while it maintains top performance on FakeAVCeleb and LAV-DF. The main contributions of this paper are:
  1. 1.
    A novel audio-visual deepfake detection methodology leveraging speech representations from both modalities to identify local cross-modal incongruities, called DiMoDif.
  2. 2.
    An extensive evaluation analysis of ablation, comparative, robustness, generalization, and in-the-wild experiments demonstrating DiMoDif’s effectiveness on deepfake detection and localization.
  3. 3.
    An inherent capability for interpreting DFD outputs through DiMoDif’s cross-modal representation incongruity.
@@ 2 Related Work
2.1 Deepfake Detection Approaches
Deepfake detection approaches considering visual-only manipulations [65, 60, 82, 40, 1, 18, 57, 66, 41], while valuable, overlook the audio component of videos and its interplay with the visual content. We compare DiMoDif with a representative set of these approaches following previous practices.
Recent deepfake detection approaches employ multimodal techniques like fusion, contrastive learning, and self-supervised learning to identify discrepancies between visual and audio streams. While these methods effectively capture audio-visual inconsistencies, they neglect the speech component of videos. For instance, attention-based frameworks analyze synchronization patterns [84], spatio-temporal models assess inter- and intra-modal disharmonies [75], and forgery-aware adaptation is applied to pre-trained ViTs [52]. Similarly, [58, 32, 86, 72] propose joint representation learning frameworks that consider inter- and intra-modal encoding mechanisms. UMMAFormer [80] adopts feature reconstruction and cross-reconstruction attention combined with an enhanced feature pyramid network, while BA-TFD & BA-TFD+ [10, 9] adopt 3DCNN and multi-scale vision Transformer backbones guided by contrastive, frame classification, and boundary matching objectives. Contrastive learning approaches compare inter-modal representations within a video [15, 14] or between videos [81]. [27] identifies audio-visual inconsistency by focusing on feature similarity, and [61] by comparing the video’s lip movements with audio-derived lip sequences. Self-supervised learning addresses the generalization issue of deepfake detection by exploiting natural audio-visual synchronization in real videos [79, 26, 29, 23, 19, 54, 77, 22]. Representation learning has also been tailored to modality emotion consistency [48] and articulatory/lip movement consistency [73]. Finally, works [44, 51, 64, 78, 5] dealing with Temporal Action Localization (TAL) are considered relevant by the deepfake detection literature.
Related works are presented in [7, 42]. These rely on a simple threshold on the divergence between VSR and ASR model outputs. However, as shown in fig. 2 this simple technique is insufficient on benchmarks with harder samples than FakeAVCeleb, such as LAV-DF and AV-Deepfake1M. In contrast, our approach extracts VSR and ASR content representations that are then processed by a Transformer-based fusion mechanism capable of identifying discriminative speech features on top of speech divergence ones. Pre-trained VSR models have been utilized as feature extractors by Haliassos et al. [30] for visual-only deepfake detection but neither audio speech features nor cross-modal incongruities were considered. Finally, [59] sets a threshold on cross-modal feature similarity deriving from an audio-visual speech self-supervised representation learning model.
@@ 2.2 Deepfake Detection Datasets
During the past years, several datasets have been proposed for video deepfake detection. Most of them focus solely on visual manipulations, thus including only the visual modality in their samples, e.g., DF-TIMIT [37], FaceForensics++ [60], DeeperForensics [33], Celeb-DF [43], WildDeepfake [85], and DF-Platter [49].
Our focus is on the detection of audio-visual deepfakes, which require content, manipulations and annotations in both modalities. Table 1 presents deepfake detection datasets that provide both visual and audio streams. DFDC [21] is a seminal large-scale benchmark in the field that contains approximately 128K samples. Although it provides audio-visual content and contains manipulations on both modalities, (i) it does not provide separate manipulation annotations per modality, (ii) fake audio parts appear in only 4% of the total number of videos, while fake visual parts in 84% [^3] Based on a later estimation: https://www.kaggle.com/datasets/basharallabadi/dfdc-video-audio-labels., and (iii) it contains videos where the depicted individual does not talk or the speaking person is not depicted and the depicted one does not talk. These features render DFDC inappropriate for audio-visual deepfake detection. KoDF [38] is another large-scale benchmark containing 237K videos of Korean subjects. Although it provides both the visual and audio components of its samples, it contains only visual manipulations, which makes it unsuitable for our problem. FakeAVCeleb [36], contains 21K manipulated and 0.5K real videos with manipulations on both modalities and corresponding annotations. Due to its size it is commonly used as an evaluation benchmark; here, we consider it for training and evaluation purposes with real video additions from the VoxCeleb2 dataset [16]. LAV-DF [10] is a recently introduced dataset that contains 136K real and fake videos with manipulations and targets on both modalities for two tasks, Deepfake Detection (DFD) and Temporal Forgery Localization (TFL). Finally, AV-Deepfake1M [8] is the largest audio-visual deepfake detection benchmark to date containing over 1.1M real and fake videos with manipulations and targets on both modalities, for both the DFD and TFL tasks.
@@ 3 Methodology
In this section we elaborate on the main components of DiMoDif, namely Discourse-related Feature Extraction (DiFE), Modality-information Differentiation (MiD), classification & regression heads, and objective function. DiMoDif’s architecture is illustrated in Figure 3.
Figure 3: The DiMoDif architecture.
@@ 3.1 Problem Formulation
Consider $(\mathfrak{v},y,\mathbf{y})\in\mathfrak{D}$ where $\mathfrak{v}=\{v,a\}$ denotes a video of dataset $\mathfrak{D}$, containing the visual $v\in\mathbb{R}^{h\times w\times c\times f}$ and the audio $a\in\mathbb{R}^{s}$ signal [^4] Utterances mainly comprise monophonic signals; conversion to single-channel is straightforward if required., with $h$ denoting the height, $w$ the width, $c$ the number of channels, $f$ the number of video frames, and $s$ the number of audio samples. $y\in\{0,1\}^{2}$ denotes the Deepfake Detection (DFD) target per modality, while $\mathbf{y}=\{y^{v}_{1},\dots,y^{v}_{f},y^{a}_{1},\dots,y^{a}_{f}\}$, with $y^{m}_{\phi}=(d^{m,s}_{\phi},d^{m,e}_{\phi},a^{m}_{\phi})$, denotes the Temporal Forgery Localization (TFL) target, assigning a forgery ground truth $a^{m}_{\phi}\in\{0,1\}$ to each modality $m\in\{v,a\}$ and frame $\phi$, along with the corresponding distances $d^{m,s}_{\phi},d^{m,e}_{\phi}\in\mathbb{R}$ between frame $\phi$ and the onset $s^{m}_{\phi}$ and offset $e^{m}_{\phi}$ of a fake interval containing $\phi$. We train deep networks $\mathcal{F}$ on the DFD task $\hat{y}=\mathcal{F}(\mathfrak{v})$ and the TFL task $\hat{\mathbf{y}}=\mathcal{F}(\mathfrak{v})$ separately, while predictions $\hat{\mathbf{y}}$ are decoded by:
$\hat{a}^{m}_{\phi}>0.5\text{, }\hat{s}^{m}_{\phi}=\phi-\hat{d}^{m,s}_{\phi}\text{, }\hat{e}^{m}_{\phi}=\phi-\hat{d}^{m,e}_{\phi}$ (1)
$\hat{\cdot}$ denotes prediction while its absence ground truth.
3.2 Discourse-related Feature Extraction (DiFE)